OpenStudy (anonymous):

solve the following system of equation. enter the x-coordinate of the solution. round your answer to the nearest tenth 5x+2=22 -2x+6y=3

OpenStudy (anonymous):

ii need help ASAP

OpenStudy (anonymous):

5x+2=22 -2 -2 5x=20 x=2

OpenStudy (anonymous):

-2x+6y=3 -2(2)+6y=3 -4+6y=3 +4 +4 6y=7 y=1.1

OpenStudy (anonymous):

x=2 and y=1.2 (since you would round 1.16)